# {tocentry}[Uncaptioned image]

### v4

@@ 1 Introduction
The QSCI method has also been demonstrated on a larger scale by Robledo-Moreno et al.34, under the name sample-based quantum diagonalization (SQD) with applications to potential energy curves of the nitrogen molecule (in a cc-pVDZ basis) and the iron-sulfur cluster model systems [2Fe-2S] and [4Fe-4S], with active spaces of (30e,20o) and (54e,36o), respectively. The demonstrations in Ref. 34 are relatively “large-scale” in nature, using CI expansion with millions of determinants. The central procedure is similar to the original QSCI paper but demonstrates the use of a large number of qubits (up to 77) and handling aspects such as treating quantum samples with incorrect particle numbers. The state preparation is handled with a local unitary cluster Jastrow ansatz (LUCJ)35, 36, with parameters assigned classically based on CCSD $t_{2}$-amplitudes. The SQD approach has since become a showcase for quantum-centric supercomputing37 and has been applied in several follow-up works, including the treatment of interaction energies with the potential energy surfaces of water and methane dimers38, spin state energetics of Fe(III)-NTA 39 and methylene40, the treatment of excited states41, inclusion of environment effects through density matrix embedding theory42, in combination with implicit-solvation models43, and combined with phaseless auxiliary- field quantum Monte Carlo44.
Separately, we note that an approach for determining the ground state of the SCI using quantum computers has also been proposed45. Unlike QSCI and SQD, determinants are selected using conventional SCI heuristics, and the focus is on leveraging variational quantum circuits to find the ground-state energy of the CI Hamiltonian.
The QSCI method introduces a potentially useful approach to determinant selection via quantum sampling, but several fundamental challenges limit its practical utility in real-world applications. First of all, the method arguably requires the preparation of a good state on the quantum device. Ideally, such a state should provide a probability distribution that closely matches the exact FCI solution since one would otherwise miss important determinants or include a significant fraction of irrelevant ones. Preparing such states is difficult, and it remains unclear how such a state preparation could take place without solving the electronic Schrödinger equation to high accuracy in the first place. Independently, two groups recently proposed methods with state-preparation protocols based on time-evolved or Hamiltonian-simulation QSCI, which could perhaps provide a workaround for this problem46, 47. Later, similar developments under the SQD moniker have independently appeared48.
@@ 2 Theory
with the CI coefficients, $c_{I}$. The included determinants can be selected by excitation level out of a reference determinant (e.g., Hartree-Fock), forming the CISD, CISDT, etc. hierarchy49, which eventually terminates at the FCI expansion when $N$-tuple excitations are included. Alternatively, as is done in SCI expansions11, 12, 13, 14, a smaller set of important determinants can be selected based on classical heuristics. This approach recognizes that only a small fraction of determinants in the FCI expansion contribute significantly to the electronic wave function for most molecular systems 50.
@@ 3 Computational Details
Hartree-Fock and CASCI (10e, 22o) wave functions, as well as one- and two-electron molecular orbital (MO) integrals (cc-pVDZ51 basis) of the nitrogen molecule, were obtained using PySCF 52. We considered structures near the equilibrium geometry ($R=1.09$Å) and along the potential energy surface $R=0.7$Å to $R=3.0$Å. Additionally, we investigated the iron-sulfur cluster [2Fe-2S] (30e,20o), where the one- and two-electron integrals were taken from the FCIDUMP file published by Li and Chan53, 54. Finally, we also include results to assess the scalability of QSCI with system size, using linear lithium hydride chains, $(\mathrm{LiH})_{n}$/STO-3G, as an example, applying full-valence CAS($2n$e, $5n$o) active spaces. For this, we consider a weakly correlated regime $R_{\mathrm{Li-H}}=1.61$, $R_{\mathrm{LiH-LiH}}=1.81$, and a strongly correlated regime with $R_{\mathrm{Li-H}}=3.22$, $R_{\mathrm{LiH-LiH}}=3.62$. Throughout the paper, we adopted the notation ($N$e, $M$o) for the active spaces, where $N$ is the number of active electrons and $M$ is the number of active orbitals.
The QSCI calculations were performed by drawing determinant samples from the CI vectors from the CASCI calculations with the probability of selecting a particular determinant given as $|c_{I}|^{2}$. The number of samples corresponds to the number of times the ansatz is measured. Our approach constitutes an ideal limit of QSCI since it assumes perfect state preparation of the exact CI vector and avoids any quantum device noise (i.e., non-ideal execution of the quantum circuits). The PyCI library55 was used to diagonalize the resulting CI Hamiltonian. HCI calculations were also performed using PyCI.
To assess the impact of an approximate ansatz in the sampling routine of QSCI, we obtained state vectors of equilibrium N2 for the unitary cluster Jastrow (UCJ)35 and local UCJ (LUCJ)36 ansatz using ffsim56. For both methods, two layers were used, and the ansatz parameters were assigned using the $t_{2}$ amplitudes of a classic CCSD simulation. This mimics the state preparation procedure used in recent works of IBM Quantum employing SQD/QSCI on quantum hardware34.
@@ 4.2 Scaled probability distributions
An additional factor to consider is that preparing an FCI-quality wave function on the quantum computer is difficult in the first place. In practice, the determinants would likely be sampled from a lower-quality wave function. We include examples of such wave functions in Figure 3, exemplified by the 2-UCJ35 and heavy-hex 2-LUCJ36 wave functions, with initial parameters assigned from CCSD $t_{2}$-amplitudes. This choice of ansatz has proven popular for previous hardware demonstrations34, although we point out that not a full 2-layer LUCJ was used in these works but rather a more approximated approach, resembling an ansatz structure between one and two layers. We find that sampling from the UCJ and LUCJ wave functions gives significantly worse results than the previously discussed sampling from the CASCI wave function. This manifests both in terms of a lower number of determinants per sample and significantly worse energies, even for relatively large CI expansions. The worse energies are likely due to a different ranking of the determinant weights, as shown in Figure S6. Curiously, LUCJ has a slightly better determinant selection than UCJ, although this is coupled with a more severe difficulty in finding new determinants and has an overall worse energy than QSCI sampled from the ideal CI vector. The less-than-stellar performance of the approximate wave functions (UCJ/LUCJ) could, at least in principle, be improved by actually optimizing the ansatz rather than just assigning parameters based on CCSD $t_{2}$ amplitudes. Such optimization has been investigated to some extent in Ref. 34 using quantum simulators. However, this optimization is likely to bring the probability distribution closer to the ground-state distribution, which, as shown above, can still lead to inefficient sampling. To the best of our knowledge, implementing this optimization on quantum hardware in practice proves to be a significant challenge and would still suffer from the fundamental problems of QSCI reported above. In general, the structure and parameter initialization of UCJ/LUCJ ansätze are not yet well-understood. For example, in Ref. 44, the authors find that LUCJ circuits with completely random parameter initializations performed better than the more established approach of assigning parameters based on CCSD $t_{2}$-amplitudes. Thus, the poorer performance of LUCJ compared to exact ground-state sampling in the present study should not be taken as a definitive indication of its potential, and the effectiveness of such approximate ansätze could improve as theoretical understanding advances.
@@ 4.3 More correlated systems: Iron-sulfur clusters
Next, we consider the more challenging electronic structure of the iron-sulfur cluster [2Fe-2S] in a (30e,20o) active space. The one- and two-electron integrals for this system were obtained from Li and Chan53, 54. We note that the same system has been studied in several QSCI works34, 41. The number of determinants in this space ($240\times 10^{6}$) is small enough that a CASCI wave function can comfortably be obtained, so we will use this as the sampling wave function for the QSCI.
@@
Figure 4 shows the convergence in the energy as a function of the number of determinants included in the CI expansions for HCI and QSCI wave functions. For the HCI results, we include results for two different initial guesses. First, starting from an HF-like determinant (blue lines), which constitutes the simple yet “bad guess”. As the orbitals from Ref. 53 are local orbitals, the HF-like determinant corresponds to filling all electrons into the left side of the iron-sulfur cluster, which is clearly an unphysical configuration. In our second approach, we start from a set of $M_{s}=0$ determinants (green lines) with five spin-up electrons on one of the iron centers (Fe 3d orbitals) and five spin-down electrons on the other iron center. This corresponds to a configuration of anti-ferromagnetically coupled spins and would be a “good guess” based on a chemical understanding of the system. It also happens to be the largest weight in the CASCI expansion. The initial guess also impacts the magnitude of the “correlation energy”, defined as the energy difference between the single-determinant Hartree-Fock and the CASCI result. For the “bad guess”, one obtains a correlation energy of $-9.5$ Hartree, while for the chemically intuitive anti-ferromagnetically coupled set of determinants, one finds a smaller correlation energy of just $-0.65$ Hartree. In Figure 4, we show the HCI results and see that the HCI expansions are also sensitive to the choice of starting determinants, with chemical intuition allowing for more compact CI expansions. For those of us who lack good chemical intuition, the good news is that after including a sufficiently large number of determinants (after around $10^{5}$), the two HCI calculations attain comparable energies, even when a bad initial guess is used. With the bad HCI starting guess (blue lines), the QSCI wave function (red lines) is initially (i.e., at large energy errors) more compact than the HCI one until a cross-over point at around 10,000 determinants (at an energy error of about 0.2 Ha). Beyond this point, for the same accuracy result, the QSCI wave function always requires more determinants than the HCI expansions. The difference is even more stark when comparing QSCI with the good starting guess of HCI expansion (green lines). Here, the QSCI wave function is consistently less compact. For example, HCI manages an energy error below 0.1 Hartree with 4,260 determinants, which is more than one order of magnitude smaller than the 50,551 determinants required by QSCI for the same energy.
@@
A careful reader may have noticed that our CI expansions (both HCI and QSCI) are much smaller than those reported in previous QSCI works (see, for example, Ref. 34 and Figure S10 therein). In Ref. 34, on the exact same [2Fe-2S] system, around $100\times 10^{6}$ determinants are required to reach an energy of around $-116.60$ Hartree, more than two orders of magnitude greater than the around $0.81\times 10^{6}$ determinants required in our calculations. The reportedly large number of determinants was especially surprising to us, considering that the entire CASCI space contains only $240\times 10^{6}$ determinants. This means that in some of the larger calculations of Ref. 34, which include nearly $200\times 10^{6}$ determinants, more than 80% of all determinants have been selected for the SCI expansion. Although the sampling component is not entirely identical, the (classically) variationally optimized LUCJ circuits employed in Ref. 34 likely approximate the ground-state distribution we consider reasonably well. Thus, we believe the discrepancy could be related to a slightly suboptimal SCI solver. A cursory glance at the implementation57 reveals that the approach relies on the kernel_fixed_space selected CI subroutine from PySCF52. As mentioned in the comments in the source code of the PySCF implementation, this subroutine is “…an inefficient dialect of Selected CI using the same structure as determinant based FCI algorithm”. In particular, this implementation does not select individual determinants. Instead, a set of $\alpha$- and $\beta$-strings are selected, and every determinant formed by all combinations thereof is included in the CI expansion. Such an algorithm may include up to quadratically more determinants in the CI expansion compared to individually selecting determinants. While this partially alleviates the sampling problem (since many more determinants are included), it somewhat contradicts the core aim of selected CI expansions – namely, to obtain good wave functions with as few determinants as possible. Notably, our results challenge the claim that for [2Fe-2S], quantum sampling is more efficient than the classic HCI method34.
@@ 4.4 Generality and Sampling Considerations
So far, we have considered two specific systems and a class of scaled distributions for sampling. With these systems, we found a clear trade-off between sampling efficiency and the compactness of the resulting CI expansion. Strengthening the generality of our results, in the Supporting Information, we provide further examples of 29 additional chemical systems derived from the W4-17-MR benchmark set58 with some additional hand-picked systems (see Figures S8-S36). The conclusions drawn so far appear to be valid across a range of chemical systems, with the typical obstacle being poor sampling efficiency (similar to N2). Of interest are systems such as B2 and twisted ethene. These systems exhibit a behavior similar to the [2Fe-2S] system in the sense that there is an intermediate region where QSCI can provide more compact CI expansions than HCI initialized from the HF determinant, although this advantage is eventually lost when reaching more sizeable CI expansions, i.e., more accurate energies. A common feature of these systems is that the Hartree-Fock determinant is a poor initial guess for HCI as it has negligible weight in the ground-state CI expansion. Therefore, when HCI is initialized from the HF determinant, there is an intermediate region with the possibility of converging to excited states. However, as the threshold $\varepsilon$ is lowered, eventually, one connects to determinants important for the ground state, which leads to rapid convergence to the ground-state energy. These drawbacks can sometimes be avoided with good chemical intuition (recall [2Fe-2S]) or starting the HCI from determinants selected from a CIS calculation (twisted ethene).
@@
A further problem arises when scaling towards larger systems. As shown in Figure 6, both QSCI and HCI incur exponentially increasing computational costs with increasing system size, which comes as no surprise, since any SCI method (independent of the sampling being done classically or quantum) inherently has an exponential computational scaling with system size59. We use linear chains of lithium hydride as an example. The bond- and intermolecular distances are selected to probe either weakly (near-equilibrium) or strongly correlated (stretched) regimes. The weakly correlated regime features a probability distribution that is similar to that of the previously discussed N2, while the strongly correlated regime has a flatter probability distribution similar to [2Fe-2S]. On a positive note, we find the sampling costs from QSCI show an exponential scaling with a smaller exponent than the corresponding complete HCI calculation time. Thus, with high-enough sampling rates, QSCI could have the potential to show usefulness.

### v5

@@ 1 Introduction
The QSCI method has also been demonstrated on a larger scale by Robledo-Moreno et al.34, under the name sample-based quantum diagonalization (SQD) with applications to potential energy curves of the nitrogen molecule (in a cc-pVDZ basis) and the iron-sulfur cluster model systems [2Fe-2S] and [4Fe-4S], with active spaces of (30e,20o) and (54e,36o), respectively. The demonstrations in Ref. 34 are relatively “large-scale” in nature, using CI expansion with millions of determinants. The central procedure is similar to the original QSCI paper but demonstrates the use of a large number of qubits (up to 77) and handling aspects such as treating quantum samples with incorrect particle numbers. The state preparation is handled with a local unitary cluster Jastrow ansatz (LUCJ)35, 36, with parameters assigned classically based on CCSD $t_{2}$-amplitudes. The SQD approach has since become a showcase for quantum-centric supercomputing37 and has been applied in several follow-up works, including the treatment of interaction energies with the potential energy surfaces of water and methane dimers38, spin state energetics of Fe(III)-NTA 39 and methylene40, the treatment of excited states41, inclusion of environment effects through density matrix embedding theory42, in combination with implicit-solvation models43, and combined with phaseless auxiliary- field quantum Monte Carlo44, 45.
Separately, we note that an approach for determining the ground state of the SCI using quantum computers has also been proposed46. Unlike QSCI and SQD, determinants are selected using conventional SCI heuristics, and the focus is on leveraging variational quantum circuits to find the ground-state energy of the CI Hamiltonian.
The QSCI method introduces a potentially useful approach to determinant selection via quantum sampling, but several fundamental challenges limit its practical utility in real-world applications. First of all, the method arguably requires the preparation of a good state on the quantum device. Ideally, such a state should provide a probability distribution that closely matches the exact FCI solution since one would otherwise miss important determinants or include a significant fraction of irrelevant ones. Preparing such states is difficult, and it remains unclear how such a state preparation could take place without solving the electronic Schrödinger equation to high accuracy in the first place. Independently, two groups recently proposed methods with state-preparation protocols based on time-evolved or Hamiltonian-simulation QSCI, which could perhaps provide a workaround for this problem47, 48. Later, similar developments under the SQD moniker have independently appeared49.
@@ 2 Theory
with the CI coefficients, $c_{I}$. The included determinants can be selected by excitation level out of a reference determinant (e.g., Hartree-Fock), forming the CISD, CISDT, etc. hierarchy50, which eventually terminates at the FCI expansion when $N$-tuple excitations are included. Alternatively, as is done in SCI expansions11, 12, 13, 14, a smaller set of important determinants can be selected based on classical heuristics. This approach recognizes that only a small fraction of determinants in the FCI expansion contribute significantly to the electronic wave function for most molecular systems 51.
@@ 3 Computational Details
Hartree-Fock and CASCI (10e, 22o) wave functions, as well as one- and two-electron molecular orbital (MO) integrals (cc-pVDZ52 basis) of the nitrogen molecule, were obtained using PySCF 53. We considered structures near the equilibrium geometry ($R=1.09$Å) and along the potential energy surface $R=0.7$Å to $R=3.0$Å. Additionally, we investigated the iron-sulfur cluster [2Fe-2S] (30e,20o), where the one- and two-electron integrals were taken from the FCIDUMP file published by Li and Chan54, 55. Finally, we also include results to assess the scalability of QSCI with system size, using linear lithium hydride chains, $(\mathrm{LiH})_{n}$/STO-3G, as an example, applying full-valence CAS($2n$e, $5n$o) active spaces. For this, we consider a weakly correlated regime $R_{\mathrm{Li-H}}=1.61$, $R_{\mathrm{LiH-LiH}}=1.81$, and a strongly correlated regime with $R_{\mathrm{Li-H}}=3.22$, $R_{\mathrm{LiH-LiH}}=3.62$. Throughout the paper, we adopted the notation ($N$e, $M$o) for the active spaces, where $N$ is the number of active electrons and $M$ is the number of active orbitals.
The QSCI calculations were performed by drawing determinant samples from the CI vectors from the CASCI calculations with the probability of selecting a particular determinant given as $|c_{I}|^{2}$. The number of samples corresponds to the number of times the ansatz is measured. Our approach constitutes an ideal limit of QSCI since it assumes perfect state preparation of the exact CI vector and avoids any quantum device noise (i.e., non-ideal execution of the quantum circuits). The PyCI library56 was used to diagonalize the resulting CI Hamiltonian. HCI calculations were also performed using PyCI.
To assess the impact of an approximate ansatz in the sampling routine of QSCI, we obtained state vectors of equilibrium N2 for the unitary cluster Jastrow (UCJ)35 and local UCJ (LUCJ)36 ansatz using ffsim57. For both methods, two layers were used, and the ansatz parameters were assigned using the $t_{2}$ amplitudes of a classic CCSD simulation. This mimics the state preparation procedure used in recent works of IBM Quantum employing SQD/QSCI on quantum hardware34.
@@ 4.2 Scaled probability distributions
An additional factor to consider is that preparing an FCI-quality wave function on the quantum computer is difficult in the first place. In practice, the determinants would likely be sampled from a lower-quality wave function. We include examples of such wave functions in Figure 3, exemplified by the 2-UCJ35 and heavy-hex 2-LUCJ36 wave functions, with initial parameters assigned from CCSD $t_{2}$-amplitudes. This choice of ansatz has proven popular for previous hardware demonstrations34, although we point out that not a full 2-layer LUCJ was used in these works but rather a more approximated approach, resembling an ansatz structure between one and two layers. We find that sampling from the UCJ and LUCJ wave functions gives significantly worse results than the previously discussed sampling from the CASCI wave function. This manifests both in terms of a lower number of determinants per sample and significantly worse energies, even for relatively large CI expansions. The worse energies are likely due to a different ranking of the determinant weights, as shown in Figure S6. Curiously, LUCJ has a slightly better determinant selection than UCJ, although this is coupled with a more severe difficulty in finding new determinants and has an overall worse energy than QSCI sampled from the ideal CI vector. The less-than-stellar performance of the approximate wave functions (UCJ/LUCJ) could, at least in principle, be improved by actually optimizing the ansatz rather than just assigning parameters based on CCSD $t_{2}$ amplitudes. Such optimization has been investigated to some extent in Ref. 34 using quantum simulators. However, this optimization is likely to bring the probability distribution closer to the ground-state distribution, which, as shown above, can still lead to inefficient sampling. To the best of our knowledge, implementing this optimization on quantum hardware in practice proves to be a significant challenge and would still suffer from the fundamental problems of QSCI reported above. In general, the structure and parameter initialization of UCJ/LUCJ ansätze are not yet well-understood. For example, in Ref. 45, the authors find that LUCJ circuits with completely random parameter initializations performed better than the more established approach of assigning parameters based on CCSD $t_{2}$-amplitudes. Thus, the poorer performance of LUCJ compared to exact ground-state sampling in the present study should not be taken as a definitive indication of its potential, and the effectiveness of such approximate ansätze could improve as theoretical understanding advances.
@@ 4.3 More correlated systems: Iron-sulfur clusters
Next, we consider the more challenging electronic structure of the iron-sulfur cluster [2Fe-2S] in a (30e,20o) active space. The one- and two-electron integrals for this system were obtained from Li and Chan54, 55. We note that the same system has been studied in several QSCI works34, 41. The number of determinants in this space ($240\times 10^{6}$) is small enough that a CASCI wave function can comfortably be obtained, so we will use this as the sampling wave function for the QSCI.
@@
Figure 4 shows the convergence in the energy as a function of the number of determinants included in the CI expansions for HCI and QSCI wave functions. For the HCI results, we include results for two different initial guesses. First, starting from an HF-like determinant (blue lines), which constitutes the simple yet “bad guess”. As the orbitals from Ref. 54 are local orbitals, the HF-like determinant corresponds to filling all electrons into the left side of the iron-sulfur cluster, which is clearly an unphysical configuration. In our second approach, we start from a set of $M_{s}=0$ determinants (green lines) with five spin-up electrons on one of the iron centers (Fe 3d orbitals) and five spin-down electrons on the other iron center. This corresponds to a configuration of anti-ferromagnetically coupled spins and would be a “good guess” based on a chemical understanding of the system. It also happens to be the largest weight in the CASCI expansion. The initial guess also impacts the magnitude of the “correlation energy”, defined as the energy difference between the single-determinant Hartree-Fock and the CASCI result. For the “bad guess”, one obtains a correlation energy of $-9.5$ Hartree, while for the chemically intuitive anti-ferromagnetically coupled set of determinants, one finds a smaller correlation energy of just $-0.65$ Hartree. In Figure 4, we show the HCI results and see that the HCI expansions are also sensitive to the choice of starting determinants, with chemical intuition allowing for more compact CI expansions. For those of us who lack good chemical intuition, the good news is that after including a sufficiently large number of determinants (after around $10^{5}$), the two HCI calculations attain comparable energies, even when a bad initial guess is used. With the bad HCI starting guess (blue lines), the QSCI wave function (red lines) is initially (i.e., at large energy errors) more compact than the HCI one until a cross-over point at around 10,000 determinants (at an energy error of about 0.2 Ha). Beyond this point, for the same accuracy result, the QSCI wave function always requires more determinants than the HCI expansions. The difference is even more stark when comparing QSCI with the good starting guess of HCI expansion (green lines). Here, the QSCI wave function is consistently less compact. For example, HCI manages an energy error below 0.1 Hartree with 4,260 determinants, which is more than one order of magnitude smaller than the 50,551 determinants required by QSCI for the same energy.
@@
A careful reader may have noticed that our CI expansions (both HCI and QSCI) are much smaller than those reported in previous QSCI works (see, for example, Ref. 34 and Figure S10 therein). In Ref. 34, on the exact same [2Fe-2S] system, around $100\times 10^{6}$ determinants are required to reach an energy of around $-116.60$ Hartree, more than two orders of magnitude greater than the around $0.81\times 10^{6}$ determinants required in our calculations. The reportedly large number of determinants was especially surprising to us, considering that the entire CASCI space contains only $240\times 10^{6}$ determinants. This means that in some of the larger calculations of Ref. 34, which include nearly $200\times 10^{6}$ determinants, more than 80% of all determinants have been selected for the SCI expansion. Although the sampling component is not entirely identical, the (classically) variationally optimized LUCJ circuits employed in Ref. 34 likely approximate the ground-state distribution we consider reasonably well. Thus, we believe the discrepancy could be related to a slightly suboptimal SCI solver. A cursory glance at the implementation58 reveals that the approach relies on the kernel_fixed_space selected CI subroutine from PySCF53. As mentioned in the comments in the source code of the PySCF implementation, this subroutine is “…an inefficient dialect of Selected CI using the same structure as determinant based FCI algorithm”. In particular, this implementation does not select individual determinants. Instead, a set of $\alpha$- and $\beta$-strings are selected, and every determinant formed by all combinations thereof is included in the CI expansion. Such an algorithm may include up to quadratically more determinants in the CI expansion compared to individually selecting determinants. While this partially alleviates the sampling problem (since many more determinants are included), it somewhat contradicts the core aim of selected CI expansions – namely, to obtain good wave functions with as few determinants as possible. Notably, our results challenge the claim that for [2Fe-2S], quantum sampling is more efficient than the classic HCI method34.
@@ 4.4 Generality and Sampling Considerations
So far, we have considered two specific systems and a class of scaled distributions for sampling. With these systems, we found a clear trade-off between sampling efficiency and the compactness of the resulting CI expansion. Strengthening the generality of our results, in the Supporting Information, we provide further examples of 29 additional chemical systems derived from the W4-17-MR benchmark set59 with some additional hand-picked systems (see Figures S8-S36). The conclusions drawn so far appear to be valid across a range of chemical systems, with the typical obstacle being poor sampling efficiency (similar to N2). Of interest are systems such as B2 and twisted ethene. These systems exhibit a behavior similar to the [2Fe-2S] system in the sense that there is an intermediate region where QSCI can provide more compact CI expansions than HCI initialized from the HF determinant, although this advantage is eventually lost when reaching more sizeable CI expansions, i.e., more accurate energies. A common feature of these systems is that the Hartree-Fock determinant is a poor initial guess for HCI as it has negligible weight in the ground-state CI expansion. Therefore, when HCI is initialized from the HF determinant, there is an intermediate region with the possibility of converging to excited states. However, as the threshold $\varepsilon$ is lowered, eventually, one connects to determinants important for the ground state, which leads to rapid convergence to the ground-state energy. These drawbacks can sometimes be avoided with good chemical intuition (recall [2Fe-2S]) or starting the HCI from determinants selected from a CIS calculation (twisted ethene).
@@
A further problem arises when scaling towards larger systems. As shown in Figure 6, both QSCI and HCI incur exponentially increasing computational costs with increasing system size, which comes as no surprise, since any SCI method (independent of the sampling being done classically or quantum) inherently has an exponential computational scaling with system size60. We use linear chains of lithium hydride as an example. The bond- and intermolecular distances are selected to probe either weakly (near-equilibrium) or strongly correlated (stretched) regimes. The weakly correlated regime features a probability distribution that is similar to that of the previously discussed N2, while the strongly correlated regime has a flatter probability distribution similar to [2Fe-2S]. On a positive note, we find the sampling costs from QSCI show an exponential scaling with a smaller exponent than the corresponding complete HCI calculation time. Thus, with high-enough sampling rates, QSCI could have the potential to show usefulness.